OpenStudy (anonymous):

i will medal and fan separate the intervals in which the function f defined on R by f(x) =x^3 -6x^2+9x +4, x in R, is increasing or decreasing . (2) show that the function f defined on R by f(x)=x^3-6x^2+9x+4, for all x in R is increasing in every interval

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@mathmate @AllTehMaffs @whydoihavetosignup1 @StudyGurl14 @sleepyjess @Somy @nothingwasthesame

OpenStudy (anonymous):

@TSwizzle

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Find the derivatives and check which intervals give you negative/positive values for the derivative. If \(f'>0\), then \(f\) is increasing; if \(f'<0\), then \(f\) is decreasing.

OpenStudy (anonymous):

ok.but how can i put it

OpenStudy (anonymous):

i mean the negative and the positive . like \[(\infty,-\infty)\]

OpenStudy (anonymous):

please just help me solve it. am lost. i know that the derivative is 3x^2-12x+9

OpenStudy (anonymous):

what should i do next

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Find the roots of the polynomial. \[3x^2-12x+9=0~~\implies~~x^2-4x+3=0~~\implies~~(x-?)(x-?)=0\]

OpenStudy (anonymous):

ok. i got that. please continue

OpenStudy (anonymous):

i.s x1=3 and x2= 1

OpenStudy (anonymous):

please what next?

OpenStudy (anonymous):

@Mendicant_Bias

OpenStudy (mendicant_bias):

Yes, those roots are correct, if I read right.

OpenStudy (anonymous):

please finish it off

OpenStudy (anonymous):

i want to really learn this and need no mistakes